How to solve this problem..the sum of 2 numbers is 63, and one is 27 larger than the other. What are the 2 numbers ____ ____?
Answer by checkley71(8403)

x+y=63 or x=63-y
y=x+27 now replace y in the first equation with (x+27) & solve for x
x=63-(x=27)
x+x=63-27
2x=36
x=36/2
x=18 answer for the smaller number.
18+27=45 the larger number.
proof
18+45=63
63=63